a merchant hass coffee worth $20.00 a pound that she wishes to mix with 30 pounds of coffee worth $90.00 a pound to get a mixture that can be sold for $40.00 a pound. How many pounds of the $20.00 coffee should be used?
Answer by mananth(16946)

--------price ---------------- quantity
Coffee I 90 ---------------- 30 pounds
coffee II 20 ---------------- x pounds
Mixture 40 ---------------- 30 + x pounds

90*30+20x=40(30 +x)
2700+20x=1200+40x
20x+-40 x=1200 - -2700
-20 x = -1500
/ -20
x = 75
75 pounds of $20 coffee II